Hi,
I am having trouble finding the right equation to use in solving the following word problem from my GRE math test prep worksheet. Can you please advise me on how to start?
Two cars started from the same point and traveled on a straight course in opposite directions for exactly 2 hours, at which time they were 208 miles apart. If one car traveled, on average, 8 mph faster than the other car, what was the average speed for each car for the 2-hour trip?
Thanks,
Julie

Answer by nerdybill(7384)

Two cars started from the same point and traveled on a straight course in opposite directions for exactly 2 hours, at which time they were 208 miles apart. If one car traveled, on average, 8 mph faster than the other car, what was the average speed for each car for the 2-hour trip?
.
Let x = speed of one car
then because "one car traveled, on average, 8 mph faster than the other car"
x+8 = speed of second car
.
Applying "distance formula":
d = rt
where
d is distance
r is rate or speed
t is time
.
"distance traveled by car one" + "distance traveled by car two" = 208
2x + 2(x+8) = 208
2x + 2x + 16 = 208
4x + 16 = 208
4x = 192
x = 48 mph (speed of one car)
.
Second car:
x+8 = 48+8 = 56 mph (speed of second car)